If your average is a 88 what's the lowest grade you can make on a test to keep your average a b?
Answer by Theo(13342)

if you assume a B average is 80, and if the weighting for all the tests is equal and you are only talking about 2 tests, then:

(x + 88)/2 = 80

multiply both sides of this equation by 2 to get:

x + 88 = 160

subtract 88 from both sides of this equation to get:

x = 160-88 = 72

lowest score you can get is 72 to keep an 80 average.

72 + 88 = 160/2 = 80